OpenStudy (anonymous):

HOW can i solve this? f(x) = x^2 + 3 g(x) = x- 7 (g o f) (2) PLEASE SHOW THE SOLUTION. THANKS! >:D<

jhonyy9 (jhonyy9):

x*2 +3 =0 --- x*2 = -3 --- x_1,_2 = +/- i radical 3 = +/- 1,73i x - 7 = 0 --- x = 7